Alpha Company signed a $200,000 mortgage on July 1, 2018 for the purchase of their new garage building. The mortgage entailed equal monthly payments of $2,600 at the end of each month. The interest rate is 5.0% per year.  How much interest expense will be paid on August 31, 2018? (Duration of mortgage was never given)
Answer #1

Interest expense july 31 = 200000*5%*1/12 = 833

Principal balance = 2600-833 = 1767

Interest expense will be paid on august 31,2018 = (200000-1767)*5%*1/12 = $826

So answer is $826